OpenStudy (anonymous):

Minimizing Surface Area You want to design an open box (no lid) with a square base that has a volume of 5000 cubic inches. (a) Express the surface area S of the box as a function of x. (b) What must the height of the box be if each side of the square base is 10 inches?

OpenStudy (anonymous):

min 2(x+y)z+xy=S st: xyz=5000=V so S=10000(x+y)/xy S'x=0 and S'y=0 x=y, z=5000/x2 So S=20000/x min S=20000/x when x=y=10 z=50